This question deals with solving the quadratic equation by completing the square. Here is the question: x^2=2x-17 Not exactly sure what do to to complete the square. Any help is much appreciated. Thanks
Answer by scott8148(6628)

x terms on one side ___ x^2 - 2x = -17

adding half of the 1st order coefficient, squared ___ [(-2) / 2]^2

x^2 - 2x + 1 = -16 ___ (x - 1)^2 = -16

taking square root ___ x - 1 = ±4i ___ x = 1 ± 4i